### v1

Abstract
Transmitting an entangled state over an extended distance is crucial for the development of quantum networks. Previous demonstrations of transmitting entangled photons over long distance using satellites or fibers have use entangled photon pairs generated from bulk crystal arrangements. An alternative approach would be to generate photon pairs using silicon-on-insulator (SOI) chips. Despite numerous proof-of-concept studies, no long range distribution has been achieved using this platform because of the challenge of getting sufficient off-chip brightness. We report a SOI platform that provides an off-chip entangled photon pair brightness of between 8,000 to 460,000 pairs per second. This exceeds previous reports by three orders of magnitude in brightness. The entanglement fidelity is 99.85(6)% and 97.90(3)% respectively. Measuring one photon locally, and transmitting the other over 93 km of deployed fiber (link loss of 40 dB), achieves a count rate of 132 pairs per second with an entanglement fidelity of 93.3(3)%, after solving the additional challenges of chromatic dispersion. The source can be pumped harder to enable transmission of entangled photons over 155 km of deployed fiber (link loss of 66 dB) at a rate of 0.7 pairs per second, with an entanglement fidelity of 87.6(5)%. These results demonstrate that SOI nanophotonic chips can perform competitively with bulk crystal sources and represent an important step toward building quantum networks using integrated nanophotonic platforms.
keywords:
Entanglement distribution, Silicon nanophotonic chip, Quantum entanglement1 Introduction
Entanglement distribution will be important for building quantum networks that provide secure communications or novel services such as blind quantum computing. Impressive distances have been achieved using satellite-based entanglement technology [1, 2, 3, 4], or with terrestrial free-space links using telescopes [5, 6]. For dense urban environments, however, optical fibers provide an alternative channel that is not limited by atmospheric conditions. Fiber-based polarization entanglement has been successfully distributed over distances up to 248 km [7, 8, 9].
A common feature of the experiments above is that the entangled photon pairs have been generated using nonlinear crystals such as bulk PPLN [7, 8, 9] or PPKTP [10]. AlGaAs chip-based sources have been explored for a 50 km fiber spool [11]. Long distance polarization entanglement distribution using photon pairs obtained from CMOS-compatible silicon nanophotonic chips have not yet been demonstrated. To overcome the link loss over long fiber stretches, the photon pair sources need to be very bright and to-date silicon photonic chips have not provided high off-chip brightness due to coupling losses [12, 13, 14, 15]. In contrast, bulk optical systems are several orders of magnitude brighter than their chip-based counterparts.
We report a bright entangled photon pair source based on a silicon nanophotonic chip that can distribute entanglement through approximately 66 dB of loss, corresponds to 155 km of deployed optical fiber in a metropolitan network. By using a novel edge coupler [16], the silicon chip is placed in a Sagnac configuration and enables direct observation of entangled photon pairs at a rate of 460,000 pairs/second. This off-chip brightness is three orders of magnitude higher than previous reports [17, 18], enabling long-distance entanglement distribution.
The entangled photons experience significant chromatic dispersion over the long distance deployed fiber route. This is compensated by using a nonlocal dispersion compensation technique [9, 19, 20] that results in an improvement in the coincidence-to-accidentals ratio (CAR) and two-photon interference theoretical fidelity by 32%. This result demonstrates that the CMOS chip sources have competitive performance when compared to sources using bulk nonlinear optics.
2 Results
2.1 On-chip entangled photon generation
The experimental setup for generating polarization entangled photons is illustrated in Fig. 1.(a, b). The test utilizes a nanophotonic chip that is an 8 mm long silicon waveguide, equipped with edge couplers at both ends, each exhibiting a loss of 0.64 dB. The silicon chip is integrated into a Sagnac fiber loop consisting of a PBS and two FPC on each arm. A continuous-wave (CW) laser with a wavelength of 1550.12 nm first passes through a standard dense wavelength division multiplexer (DWDM) filter and is then evenly split by the PBS. The vertically polarized (V) beam is converted to horizontal polarization (H) with fiber polarization controllers (FPC) before entering the chip. A short section of small-core fiber (UHNA7), less than 1 cm in length, is spliced to the tip of coupling fiber to assist mode conversion with a splicing loss of 0.03 dB.
A pump beam injected into the chip causes spontaneous four wave mixing (SFWM) to occur. A pair of pump photons at frequency are annihilated, generating a new photon pair, following the energy conservation equation . [21] To create entangled photon pairs, the pump is injected from both ends of the waveguide simultaenously, leading to the generation of photon pairs propagating in both directions. These pairs can be collected from either side of the waveguide using the same fiber as pump laser. One pair is horizontally (H) polarized and the other pair is converted to vertically (V) polarized by FPC. As the H and V state photons travel through the same fiber, no phase shift occurs between the two arms. By superposing the H and V polarized photon pairs from both paths at the PBS, the maximally entangled Bell state is constructed at the output port of the PBS:
The entangled photons are reflected by a DWDM module centered at 1550.12 nm to the reflection arm. A narrow band portion (0.8 nm) of photon pairs are then separated using 100 Ghz DWDM modules centered at 1561.01 nm (idler) and 1539.37 nm (signal). The broadband nature of the entangled photons generated from the SFWM process in the straight Si waveguide allows for the cascading of additional filters to multiplex more channels of entangled photons. [14, 11]
2.2 Entangled source characterization
Once the signal and idler photons are filtered, they first pass through two quarter-wave plates (QWP) to compensate for polarization distortion caused by the connecting fibers. Each arm of the entangled photons are then projected into different polarization bases, using half-wave plates (HWP) and polarizing beam splitters. During source characterization, a short length of fiber replaces the deployed fibers, directly connecting the source to the measurement setup.
The HWP in the signal arm is fixed at , while the HWP in the idler arm rotates for . Photons are detected by SNSPDs, with the coincidence count recorded during the rotation. The single photons are detected with SNSPDs, with the efficiency for idler channel at 45% while efficiency for signal channel is at 49%. Fig. 2.(a) displays the observed raw coincidence rate as a function of .
Source type | Detector (eff.) | Brightness (cps) | Fidelity | Ref |
Si waveguide | APD (10%) | 4 | 94.8(41)% | [13] |
Si waveguide | APD (20%) | 140 | 97.7(8)% | [14] |
Si waveguide | SNSPD (85%) | 360 | 97.3(16)% | [18] |
Si ring | SPD (10%) | 6 | 96.3(15)% | [12] |
Si ring | SNSPD (81%) | 60 | 94(2)% | [22] |
SiN ring | SNSPD (60%) | 400 | 90.0(-)% | [17] |
[…]
SNSPD (45%, 49%) | 30,000 | 99.80(4)% | ||
SNSPD (45%, 49%) | 280,000 | 98.82(3)% | ||
SNSPD (45%, 49%) | 460,000 | 97.90(3)% | ||
SNSPD (56%) | 2,800,000* | 93.9(7)% |
*This is measured with a 15 dB attenuator inserted in both the signal and idler arms before the SNSPDs to avoid detector saturation. The decreasing fidelity with increasing brightness is due to the number of multiple photon pairs present within the coincidence time window of this system. This effect can be further minimized if the system timing jitter and coincidence window is reduced.
The maximum observed coincidence at the peak of the coincidence curves is 230 kcps, corresponding to an overall brightness of 460,000 entangled photon pairs a second. The fidelity to the bell state is 97.90(3)%. This source is over 3 orders of magnitude brighter than sources in other literature reports. This comparison is provided in Fig. 3. When the source is operating at low brightness conditions, fidelity can be enhanced to 99.85(6)%, and the experimental data detailed in Table 1.
2.3 Asymmetric entanglement distribution over 93 km distance
The signal photons are first connected to the NUS-SUTD-NUS deployed fiber which is 93 km, exhibiting a 35.4 dB loss as measured by an optical time domain reflectometer (OTDR). The idler photons are detected at the source. The propagation loss is calculated to be 0.38 dB/km, which is higher than the typical 0.16 dB/km observed in fiber spools or 0.25 dB/km in submarine fibers [7, 8]. This increased loss is primarily due to the connection junctions in practical metropolitan fibers. When the dispersion compensation module (DCM) is added, 4.3 dB extra loss is also included to the fiber system.
To evaluate the entanglement quality after transmission through the NUS-SUTD-NUS fiber loop, we perform two-fold coincidence measurements equivalent to local testing. The idler arm’s polarization analysis setup is set to H, V, D, and A bases. The signal arm’s HWP is rotated through to measure the visibility. As shown in Fig. 2.(b), observed raw count rate of entangled photons is 132 cps, while the fidelity is 93.3(3)% in the 10 second measurement duration. The high fidelity attributed to effective background count filtering and nonlocal dispersion compensation.
Source type | Fiber type | Fiber loss | DOF | Rate (cps) | Fidelity | Ref |
Sagnac PPLN | Submarine (96 km) | 22 dB | pol. | 366 | 95.3(3)% | [7] |
Sagnac PPLN | Submarine (192 km) | 48 dB | pol. | 8.6 | 95(1)% | [8] |
Sagnac PPLN | Metropolitan (248 km) | 71.5 dB | pol. | 9 | 93(-)% | [9] |
Sagnac PPLN | Ultra low loss spool (404 km) | 87 dB | pol. | 0.03 | 91.3(-)% | [23] |
[…]
Metropolitan (155 km) | 66 dB | 0.7 | 87.6(5)% | |||
[…]
2.4 Symmetric entanglement distribution over 155 km deployed fiber
When only the signal photons are deployed, the timing delay between the signal and idler arms are observed to experience a cyclic drift. This cyclic drift is attributed to diurnal temperature fluctuations that cause the length of the fibers to expand and contract. This effect is reduced when both signal and idler photons are deployed over similar lengths of fiber. This “symmetric" experiment was performed by connecting the idler photons to a fiber loop that connected NUS and NTU. In this setup, the overall length traversed by the photon pair exceeds 155 km, and the overall system loss reached 66 dB (56 dB from fiber, 10 dB from DCM). By increasing the pump power, the entangled photon pair rate after transmission was 0.7 cps with a fidelity of 87.6(5)%. Accounting for fiber loss, the off-chip photon pair rate was 2.8 Mcps and the estimated fidelity was 93.9(7)% as shown in Fig. 3. The degradation is primarily due to the effects of polarization mode dispersion (PMD). [9, 24, 25, 26, 27] A comparison of the results from this work with similar work in the literature regarding fiber-based entanglement distribution is provided in Table 2.
3 Conclusion
In this report we have demonstrated a bright entangled photon pair source based on a silicon nanophotonic chip. The source was sufficiently bright that entanglement could be distributed over 66 dB of loss (in our experiment this is equivalent to 155 km of deployed fiber and a dispersion compensation module). The observed brightness and fidelity of the nanophotonic chip source is competitive when compared against work based on bulk crystal sources. [7, 8, 9] This work marks a milestone in the capability of silicon nanophotonic chips for enabling long range quantum networks. This source is also useful for satellite to ground experiments because it exceeds the loss of the Micius double-downlink where the link loss was 62 dB. [2, 28].
There are a number of avenues for further integration and improvement. For example, many of the filters are based on fiber components and further integration [29, 30, 31] is possible in this aspect enhancing the stability and scalability of the setup. We note also that the nonlocal dispersion compensation is performed using a dispersion compensated module that has a significant transmission loss, which should be reduced with further engineering [32, 33]. We note also that there is a huge potential for utilising dense-wavelength division multiplexing with the silicon nanophotonic chip. This will expand the number of channels accessible. [23, 34, 22] Finally, other degrees of freedom, such as the time-bins could be added to the setup. [35, 36, 37] Taken together, these modifications will dramatically increase the capacity for quantum information transfer within the network.
4 Methods
4.1 Chip design and coupling strategy
The chip involved in the experiment is a 8 mm long silicon waveguide with cross sectional dimension 650 nm in width and 250 nm in height. The whole chip is covered by thick SiO2 cladding layer. On the chip edges, two-stage inverse tapers are fabricated to expand the chip mode to reduce the mode mismatch with that of the UHNA7 fiber. The L0 section acts as buffer region to protect the edge coupler during dicing process, while the length of L1 and L2 tapers are and respectively. Short section of UHNA7 fiber (less than 1 cm) is spliced to SMF-28 fiber to assist light collection. The loss is optimized to 0.03 dB per splicing junction by performing an in-situ monitoring of the loss during splicing process. [16] The measured coupling loss is 0.64 dB per facet and has nice long term stability of dB for more than ten days without using any adhesive.
4.2 Photon pair wavelength selection
In the SFWM process, two pump photons are absorbed simultaneously, elevating the system to a virtual energy level, followed by the emission of a signal and idler photon to conserve both energy and momentum. The efficiency of this nonlinear interaction is highly dependent on the phase matching conditions. The coincidence-to-accidentals ratio (CAR) quantifies the ratio of true photon coincidences (from genuine photon pairs) to accidental coincidences, such as those arising from multi-photon emissions or system noise. A high CAR is essential for ensuring high entanglement fidelity, as it directly correlates with the purity of the detected entangled state by reducing the likelihood of accidental events degrading the measurement. Therefore, maximizing CAR requires lower multi-photon events and system noise, such as residual pump photons and Raman noise.
To select the optimal wavelengths for the highest CAR and entangled photon quality, a 1550.12 nm CW laser is employed to unidirectional pumping for the same chip used in the entanglement generation experiment. The generated photon pairs, signal and idler, were spectrally broad due to the nature of phase matching in silicon. Two 0.8 nm bandwidth filters with central wavelength symmetrical about the pump are installed to scan the pair generation rate across different wavelength channels.
As illustrated in Fig. 5.(a), the raw photon coincidence rate was highest when the signal and idler wavelengths were closest to the pump wavelength, indicating that phase matching was most efficient at shorter detunings from the pump. However, this increased photon pair generation does not directly translate into a higher CAR, as shown in Fig. 5.(b), where the CAR peaks approximately 10 nm away from the pump wavelength. This discrepancy can be explained by the balance between photon pair generation and system noise. While photon pairs are generated more efficiently near the pump wavelength, the proximity of the signal and idler to the pump results in less effective rejection of residual pump photons as shown in Fig. 5.(c). These residual pump photons contribute to accidental coincidences, lowering the CAR. On the other hand, when the signal and idler photons are too far from the pump, the SFWM process itself becomes less efficient, reducing the photon pair generation rate and thereby lowering the CAR due to the lower coincidence rate. The observed peak in CAR at a 10 nm separation from the pump wavelength results from the optimal trade-off between reducing system noise and maintaining sufficient photon pair generation.
4.3 Methods to improve the fidelity
In this experiment the deployed fiber is not dark and contains broadband noise within the optical channel of interest. This noise lowers the CAR value affecting the fidelity of the observed quantum state. The relationship between fidelity and CAR is given by . A 0.8 nm filter is installed before the SNSPDs, and the dark count noises are effectively removed from around 4 Mcps to less than 200 cps. The CAR is increased from 2 to 15, which corresponds to a theoretical fidelity improvement from 67% to 94%.
Additionally, chromatic dispersion presents a significant challenge to achieving high entanglement fidelity. The dispersion value at 1539 nm is around ps/(nmkm) in the fiber. Since the bandwidth of signal photons is 0.8 nm while the fiber distance is 93 km, the dispersion caused by the deployed fiber is 1339 ps. In comparison, the overall timing jitter of the SNSPDs including the time tagger, is approximately 80 ps. As a result, measuring the photon pair rate after signal photons transmitted through the long fiber requires a larger coincidence window around 1 ns. This larger coincidence window increases accidental counts, greatly reducing the CAR value. To solve this, a dispersion compensation module (DCM) is inserted in the idler arm. The total dispersion compensation provided by the DCM at 1561 nm is around -1360 ps/nm, as data provided by the manufacturer. Effective dispersion compensation raises the CAR from 15 to 91, corresponding to an increase in theoretical fidelity from 94% to 99%. However, the measured fidelity in practice is lower than this theoretical value, primarily due to the effects of PMD. It is important to note that the DCM introduces 4.3 dB insertion loss, which does impact the coincidence rate. This DCM-related loss may be mitigated in future work where the waveguide platform could be designed to generate photons with narrower wavelength bandwidth. [38, 12, 22, 17, 39]
4.4 System calibration
The polarization shift caused by birefringence is compensated by the QWP and HWP in the measurement setup. To calibrate the system once the NUS-NTU-NUS quantum channel is established, we disconnect the H polarization arm of Sagnac loop and inject a 13 dBm laser beam at 1561.01 nm into the setup. This wavelength matches that of the idler photons, allowing the laser to go through the same route as idler photons do. After transmission through the PBS in the measurement setup, the laser is detected by a free space powermeter. The QWP and HWP are then adjusted to minimize the detected power, with the final HWP angle designated as the V state for the idler arm. The output power difference between HWP’s H and V position is 37 dB which is limited by the extinction ratio of the polarization components. A similar procedure is employed for the signal arm, using a laser at 1539.37 nm. Once the waveplates’ angles are calibrated, it is crucial that the QWPs remain stationary during the two-photon interference measurement to ensure accuracy. In the local measurement case, once the photon pairs are collected with the UHNA7 fiber, the insertion loss from Sagnac loop to the fiber connector before SNSPD is 2.74 dB for signal arm and 2.72 dB for idler arm. This includes all the fiber based and free space optical components in the experiment.
4.5 Single photon detection and coincidence counting
In the 93 km entanglement distribution experiment, the single photons are detected with SNSPDs, with the efficiency for idler channel at 45% while efficiency for signal channel is at 78%. Single photon count rate is 4.8 kcps for signal arm and 5.6 Mcps for idler arm. The delay cause by the deployed fiber is 457,369,970 ps which is equal to the distance measured by OTDR. The coincidence window used in local measurement is 200 ps while with deployed fiber is 60 ps for accidental suppression and higher CAR. The narrow does result in a lower photon pair rate. Reducing the background noise (detector dark counts, Raman noise in the source, and stray light coupling into the fiber network) should further improve the CAR and fidelity. When the deployed fiber distance is extended to 155 km, the singles count rate on signal and idler arm are around 15 kcps and 40 kcps respectively. Although this low single count rate will not saturate the detector, the SNSPD are set to 56% efficiency to reduce the dark count to 10 cps. The optical delay caused by the 62 km fiber is 302,113,173 ps which is equal to the length characterization by OTDR.
4.6 Fiber link information
A measurement of polarization stability of deployed fibers over a continuous period of 8 hours is conducted and the polarization state rotates (93 km link) and (62 km link) on the Poincare sphere. The results in Fig. 7. (a, b) indicate a relatively stable polarization behavior.
Additionally, signal crosstalk from other channels in metropolitan fiber networks can introduce background noise, which also affects the quantum state fidelity. The NUS-SUTD-NUS (or NUS-NTU-NUS) fiber loop experiences approximately 4 Mcps (2 kcps) of background noise without spectral filtering. This noise is broadband and can be effectively mitigated using a 0.8 nm bandwidth DWDM filter. After inserting the filter, the background noise in both channels is reduced to less than 200 cps.
Acknowledgements
This research is supported by the National Research Foundation, Singapore, and A*STAR under its CQT Bridging Grant and Quantum Engineering Programme (Award No. NRF2021-QEP2-01-P02), QEP2 grant: NRF2021-QEP2-01-P01, and US Air Force grant (FA2386-21-1-0076). D. Tan acknowledges funding from the National Research Foundation, Singapore, and A*STAR under its QEP 2.0 Programme (Award No. NRF2022-QEP2-01-P08), the National Research Foundation Investigatorship (NRF-NRFI08-2022-0003) and the Ministry of Education Tier 2 Grant (Award No. MOE2019-T2-2-178). We acknowledge funding support from the National Research Foundation, Singapore and A*STAR under its Quantum Engineering Programme (National Quantum-Safe Network, NRF2021-QEP2-04-P01) and Netlink Trust for the provisioning of the fiber network. The authors acknowledge the assistance of setup alignment and literature review from Arya Chowdhury, and Lim En Teng, and consultation from Dr. Tanvirul Islam.
Declarations
The authors declare no conflicts of interest.

### v2

@@
-
-Abstract
-Transmitting an entangled state over an extended distance is crucial for the development of quantum networks.
-Previous demonstrations of transmitting entangled photons over long distance using satellites or fibers have use entangled photon pairs generated from bulk crystal arrangements.
-An alternative approach would be to generate photon pairs using silicon-on-insulator (SOI) chips.
-Despite numerous proof-of-concept studies, no long range distribution has been achieved using this platform because of the challenge of getting sufficient off-chip brightness.
-We report a SOI platform that provides an off-chip entangled photon pair brightness of between 8,000 to 460,000 pairs per second.
-This exceeds previous reports by three orders of magnitude in brightness.
+Entanglement distribution over 155 km metropolitan fiber using a CMOS-compatible silicon chip SUMMARY
+Entanglement distribution over 155 km metropolitan fiber using a CMOS-compatible silicon chip
+SUMMARY
+Transmitting entangled states over long distances is crucial for developing quantum networks.
+Previous demonstrations using satellites or fibers relied on photon pairs generated from bulk crystal arrangements.
+Polarization entanglement distribution based on CMOS-compatible silicon chips has long been restricted to lab-scale demonstrations spanning only a few meters, due to the difficulty of achieving sufficient off-chip brightness.
+We report a silicon chip platform that provides an off-chip entangled photon pair brightness ranging from 8,000 to 460,000 pairs per second, exceeding previous reports by three orders of magnitude.
+The entanglement fidelity reaches 99.85(6)% and 97.90(3)%, respectively.
+After addressing key challenges in long-distance entanglement distribution over deployed fiber, including phase drift and chromatic dispersion, entangled photons were successfully distributed over 155 km (66 dB loss).
+These results demonstrate that CMOS-compatible silicon chips can perform competitively with bulk crystal sources and represent an important step toward scalable, chip-based quantum networks.
+KEYWORDS
+Entanglement distribution, CMOS-compatible silicon chip, Quantum network
+INTRODUCTION
+Entanglement distribution will be important for building quantum networks that provide secure communications or novel services such as blind quantum computing.
+Impressive distances have been achieved using satellite-based entanglement technology 1, 2, 3, 4, or with terrestrial free-space links using telescopes 5, 6.
+For dense urban environments, however, optical fibers provide an alternative channel that is not limited by atmospheric conditions.
+A common strategy for fiber-based transmission is to use time-bin entanglement 7, 8, 9, 10, 11, which is robust against polarization drift in aerial fibers 12.
+However, time-bin entanglement typically relies on unbalanced interferometers which demands precise phase stabilization, and structurally leads to a 3 dB measurement loss.
+In the stable underground fiber networks, polarization entanglement offers a more direct and flexible platform for state preparation and analysis.
+It is also inherently compatible with satellite-to-ground free-space links and supports hybrid architectures that span fiber and free-space channels 13, 6, 14, 15.
+The primary challenge for polarization-entangled photons is polarization mode dispersion (PMD) in fiber, but effective mitigation strategies have recently been proposed 16, 17.
+Bulk nonlinear crystal-based sources such as PPLN 18, 19, 20 and PPKTP 21 have successfully generated bright polarization-entangled photons, which have been distributed over distances up to 248 km using deployed fiber 18, 19, 20 and 404 km using fiber spools 22.
+In contrast, polarization-entangled photons from CMOS-compatible silicon nanophotonic chips have long been limited to lab-scale demonstrations over just a few meters 23, 24.
+To overcome the link loss over long fiber stretches, the photon pair sources need to be very bright and to-date silicon photonic chips have not provided high off-chip brightness due to coupling losses 25, 26, 27, 28, 29, whereas bulk optical systems are several orders of magnitude brighter than their chip-based counterparts.
+In this work, we present a system-level optimization approach that addresses several challenges.
+This enables silicon photonic chips to operate in a regime previously dominated by bulk-crystal systems.
+We report a bright polarization entangled photon pair source based on a silicon nanophotonic chip that can distribute entanglement through approximately 66 dB of loss, corresponds to 155 km of deployed optical fiber in a metropolitan network.
+By using a novel edge coupler, the silicon chip is placed in a Sagnac configuration and enables direct observation of entangled photon pairs at a rate of between 8,000 to 460,000 pairs per second.
+This off-chip brightness is three orders of magnitude higher prior silicon-based reports 28, 27, 26, 30, 23, 24, 31, 32.
 The entanglement fidelity is 99.85(6)% and 97.90(3)% respectively.
-Measuring one photon locally, and transmitting the other over 93 km of deployed fiber (link loss of 40 dB), achieves a count rate of 132 pairs per second with an entanglement fidelity of 93.3(3)%, after solving the additional challenges of chromatic dispersion.
+Measuring one photon locally, and transmitting the other over 93 km of deployed fiber (link loss of 40 dB), achieves a count rate of 132 pairs per second with an entanglement fidelity of 93.3(3)%.
+To mitigate chromatic dispersion over long fiber distances, we implement nonlocal dispersion compensation 20, 33, 34, improving coincidence-to-accidental ratio (CAR) and fidelity upper limit by 32%.
 The source can be pumped harder to enable transmission of entangled photons over 155 km of deployed fiber (link loss of 66 dB) at a rate of 0.7 pairs per second, with an entanglement fidelity of 87.6(5)%.
-These results demonstrate that SOI nanophotonic chips can perform competitively with bulk crystal sources and represent an important step toward building quantum networks using integrated nanophotonic platforms.
-keywords:
-Entanglement distribution, Silicon nanophotonic chip, Quantum entanglement1 Introduction
-Entanglement distribution will be important for building quantum networks that provide secure communications or novel services such as blind quantum computing.
-Impressive distances have been achieved using satellite-based entanglement technology [1, 2, 3, 4], or with terrestrial free-space links using telescopes [5, 6].
-For dense urban environments, however, optical fibers provide an alternative channel that is not limited by atmospheric conditions.
-Fiber-based polarization entanglement has been successfully distributed over distances up to 248 km [7, 8, 9].
-A common feature of the experiments above is that the entangled photon pairs have been generated using nonlinear crystals such as bulk PPLN [7, 8, 9] or PPKTP [10].
-AlGaAs chip-based sources have been explored for a 50 km fiber spool [11].
-Long distance polarization entanglement distribution using photon pairs obtained from CMOS-compatible silicon nanophotonic chips have not yet been demonstrated.
-To overcome the link loss over long fiber stretches, the photon pair sources need to be very bright and to-date silicon photonic chips have not provided high off-chip brightness due to coupling losses [12, 13, 14, 15].
-In contrast, bulk optical systems are several orders of magnitude brighter than their chip-based counterparts.
-We report a bright entangled photon pair source based on a silicon nanophotonic chip that can distribute entanglement through approximately 66 dB of loss, corresponds to 155 km of deployed optical fiber in a metropolitan network.
-By using a novel edge coupler [16], the silicon chip is placed in a Sagnac configuration and enables direct observation of entangled photon pairs at a rate of 460,000 pairs/second.
-This off-chip brightness is three orders of magnitude higher than previous reports [17, 18], enabling long-distance entanglement distribution.
-The entangled photons experience significant chromatic dispersion over the long distance deployed fiber route.
-This is compensated by using a nonlocal dispersion compensation technique [9, 19, 20] that results in an improvement in the coincidence-to-accidentals ratio (CAR) and two-photon interference theoretical fidelity by 32%.
-This result demonstrates that the CMOS chip sources have competitive performance when compared to sources using bulk nonlinear optics.
-2 Results
-2.1 On-chip entangled photon generation
+This result demonstrates that the CMOS chip sources, can compete with crystal-based systems for long-range quantum networking.
+RESULTS
+On-chip entangled photon generation
 The experimental setup for generating polarization entangled photons is illustrated in Fig.
@@
 The test utilizes a nanophotonic chip that is an 8 mm long silicon waveguide, equipped with edge couplers at both ends, each exhibiting a loss of 0.64 dB.
-The silicon chip is integrated into a Sagnac fiber loop consisting of a PBS and two FPC on each arm.
+The silicon chip is integrated into a Sagnac fiber loop consisting of a PBS and two fiber polarization controllers (FPC) on each arm.
 A continuous-wave (CW) laser with a wavelength of 1550.12 nm first passes through a standard dense wavelength division multiplexer (DWDM) filter and is then evenly split by the PBS.
-The vertically polarized (V) beam is converted to horizontal polarization (H) with fiber polarization controllers (FPC) before entering the chip.
+The vertically polarized (V) beam is converted to horizontal polarization (H) with FPC before entering the chip.
 A short section of small-core fiber (UHNA7), less than 1 cm in length, is spliced to the tip of coupling fiber to assist mode conversion with a splicing loss of 0.03 dB.
 A pump beam injected into the chip causes spontaneous four wave mixing (SFWM) to occur.
-A pair of pump photons at frequency are annihilated, generating a new photon pair, following the energy conservation equation .
-[21] To create entangled photon pairs, the pump is injected from both ends of the waveguide simultaenously, leading to the generation of photon pairs propagating in both directions.
+A pair of pump photons at frequency are annihilated, generating a new photon pair, following the energy conservation equation 35.
+To create entangled photon pairs, the pump is injected from both ends of the waveguide simultaenously, leading to the generation of photon pairs propagating in both directions.
 These pairs can be collected from either side of the waveguide using the same fiber as pump laser.
@@
 A narrow band portion (0.8 nm) of photon pairs are then separated using 100 Ghz DWDM modules centered at 1561.01 nm (idler) and 1539.37 nm (signal).
-The broadband nature of the entangled photons generated from the SFWM process in the straight Si waveguide allows for the cascading of additional filters to multiplex more channels of entangled photons.
-[14, 11]
-2.2 Entangled source characterization
+The broadband nature of the entangled photons generated from the SFWM process in the straight Si waveguide allows for the cascading of additional filters to multiplex more channels of entangled photons 28, 36.
+Entangled source characterization
 Once the signal and idler photons are filtered, they first pass through two quarter-wave plates (QWP) to compensate for polarization distortion caused by the connecting fibers.
@@
 Fig.
-2.(a) displays the observed raw coincidence rate as a function of .
-Source type | Detector (eff.) | Brightness (cps) | Fidelity | Ref |
-Si waveguide | APD (10%) | 4 | 94.8(41)% | [13] |
-Si waveguide | APD (20%) | 140 | 97.7(8)% | [14] |
-Si waveguide | SNSPD (85%) | 360 | 97.3(16)% | [18] |
-Si ring | SPD (10%) | 6 | 96.3(15)% | [12] |
-Si ring | SNSPD (81%) | 60 | 94(2)% | [22] |
-SiN ring | SNSPD (60%) | 400 | 90.0(-)% | [17] |
-[…]
-SNSPD (45%, 49%) | 30,000 | 99.80(4)% | ||
-SNSPD (45%, 49%) | 280,000 | 98.82(3)% | ||
-SNSPD (45%, 49%) | 460,000 | 97.90(3)% | ||
-SNSPD (56%) | 2,800,000* | 93.9(7)% |
-*This is measured with a 15 dB attenuator inserted in both the signal and idler arms before the SNSPDs to avoid detector saturation.
-The decreasing fidelity with increasing brightness is due to the number of multiple photon pairs present within the coincidence time window of this system.
-This effect can be further minimized if the system timing jitter and coincidence window is reduced.
+2.(a) displays the observed raw coincidence rate as a function of , with the error bars estimated from Poisson statistics.
 The maximum observed coincidence at the peak of the coincidence curves is 230 kcps, corresponding to an overall brightness of 460,000 entangled photon pairs a second.
@@
 When the source is operating at low brightness conditions, fidelity can be enhanced to 99.85(6)%, and the experimental data detailed in Table 1.
-2.3 Asymmetric entanglement distribution over 93 km distance
+To further verify the presence of entanglement, a Bell test was performed by measuring the violation of the Clauser-Horne-Shimony-Holt (CHSH) inequality.
+Polarization correlations in four measurement bases were recorded, and the CHSH parameter was calculated for the source at various brightness levels.
+The highest observed value is 2.82(2) at 8 kcps brightness, and remains well above the classical bound of 2 even at the highest measured pair rates.
+| Source type | Detector (eff.) | Brightness (cps) | Fidelity | CHSH | Ref |
+| Si waveguide | APD (10%) | 4 | 94.8(41)% | - | 27 |
+| Si waveguide | APD (20%) | 140 | 97.7(8)% | 2.7(1) | 28 |
+| Si waveguide | SNSPD (85%) | 360 | 97.3(16)% | - | 32 |
+| Si ring | SPD (10%) | 6 | 96.3(15)% | - | 26 |
+| Si ring | SNSPD (81%) | 60 | 94(2)% | - | 30 |
+| SiN ring | SNSPD (60%) | 400 | 90.0(-)% | - | 31 |
+| Si waveguide | SNSPD (45%, 49%) | 8,000 | 99.85(6)% | 2.82(2) | This work |
+| SNSPD (45%, 49%) | 30,000 | 99.80(4)% | 2.815(8) | ||
+| SNSPD (45%, 49%) | 280,000 | 98.82(3)% | 2.777(3) | ||
+| SNSPD (45%, 49%) | 460,000 | 97.90(3)% | 2.729(2) | ||
+| SNSPD (56%) | 2,800,000* | 93.9(7)% | 2.44(3) |
+*This is measured with a 15 dB attenuator inserted in both the signal and idler arms before the SNSPDs to avoid detector saturation.
+The decreasing fidelity with increasing brightness is due to the number of multiple photon pairs present within the coincidence time window of this system.
+This effect can be further minimized if the system timing jitter and coincidence window is reduced.
+Asymmetric entanglement distribution over 93 km distance
 The signal photons are first connected to the NUS-SUTD-NUS deployed fiber which is 93 km, exhibiting a 35.4 dB loss as measured by an optical time domain reflectometer (OTDR).
 The idler photons are detected at the source.
-The propagation loss is calculated to be 0.38 dB/km, which is higher than the typical 0.16 dB/km observed in fiber spools or 0.25 dB/km in submarine fibers [7, 8].
+The propagation loss is calculated to be 0.38 dB/km, which is higher than the typical 0.16 dB/km observed in fiber spools or 0.25 dB/km in submarine fibers 18, 19.
 This increased loss is primarily due to the connection junctions in practical metropolitan fibers.
@@
 The high fidelity attributed to effective background count filtering and nonlocal dispersion compensation.
-Source type | Fiber type | Fiber loss | DOF | Rate (cps) | Fidelity | Ref |
-Sagnac PPLN | Submarine (96 km) | 22 dB | pol. | 366 | 95.3(3)% | [7] |
-Sagnac PPLN | Submarine (192 km) | 48 dB | pol. | 8.6 | 95(1)% | [8] |
-Sagnac PPLN | Metropolitan (248 km) | 71.5 dB | pol. | 9 | 93(-)% | [9] |
-Sagnac PPLN | Ultra low loss spool (404 km) | 87 dB | pol. | 0.03 | 91.3(-)% | [23] |
-[…]
-Metropolitan (155 km) | 66 dB | 0.7 | 87.6(5)% | |||
-[…]
-2.4 Symmetric entanglement distribution over 155 km deployed fiber
+Near-symmetric entanglement distribution over 155 km deployed fiber
 When only the signal photons are deployed, the timing delay between the signal and idler arms are observed to experience a cyclic drift.
@@
 This effect is reduced when both signal and idler photons are deployed over similar lengths of fiber.
-This “symmetric" experiment was performed by connecting the idler photons to a fiber loop that connected NUS and NTU.
-In this setup, the overall length traversed by the photon pair exceeds 155 km, and the overall system loss reached 66 dB (56 dB from fiber, 10 dB from DCM).
+This “near-symmetric” experiment was performed by connecting the idler photons to a fiber loop that connected NUS and NTU.
+In this setup, the overall length traversed by the photon pair exceeds 155 km, and the overall system loss reached 66 dB (56 dB from fiber, 10 dB from 160 km version DCM).
+Although the final coincidence detection is performed locally at NUS, both the signal and idler photons are independently transmitted over two separate long-distance fiber link before measurement.
+This effectively simulates a distributed entanglement scenario, as the entangled photons travel through metropolitan deployed fibers and experience actual losses, dispersion, and changes from the environment.
 By increasing the pump power, the entangled photon pair rate after transmission was 0.7 cps with a fidelity of 87.6(5)%.
@@
 3.
-The degradation is primarily due to the effects of polarization mode dispersion (PMD).
-[9, 24, 25, 26, 27] A comparison of the results from this work with similar work in the literature regarding fiber-based entanglement distribution is provided in Table 2.
-3 Conclusion
+The degradation is primarily due to the effects of polarization mode dispersion (PMD) 20, 38, 39, 40, 41.
+A comparison of the results from this work with similar work in the literature regarding fiber-based entanglement distribution is provided in Table 2.
+| Source type | Fiber type | Fiber loss | DOF | Rate (cps) | Fidelity | Ref |
+| Sagnac PPLN | Submarine (96 km) | 22 dB | pol. | 366 | 95.3(3)% | 18 |
+| Sagnac PPLN | Submarine (192 km) | 48 dB | pol. | 8.6 | 95(1)% | 19 |
+| Sagnac PPLN | Metropolitan (248 km) | 71.5 dB | pol. | 9 | 93(-)% | 20 |
+| Sagnac PPLN | Ultra low loss spool (404 km) | 87 dB | pol. | 0.03 | 91.3(-)% | 22 |
+| Sagnac Si chip | Metropolitan (93 km) | 39.7 dB | pol. | 132 | 93.3(3)% | This work |
+| Metropolitan (155 km) | 66 dB | 0.7 | 87.6(5)% |
+Discussion
 In this report we have demonstrated a bright entangled photon pair source based on a silicon nanophotonic chip.
-The source was sufficiently bright that entanglement could be distributed over 66 dB of loss (in our experiment this is equivalent to 155 km of deployed fiber and a dispersion compensation module).
-The observed brightness and fidelity of the nanophotonic chip source is competitive when compared against work based on bulk crystal sources.
-[7, 8, 9] This work marks a milestone in the capability of silicon nanophotonic chips for enabling long range quantum networks.
-This source is also useful for satellite to ground experiments because it exceeds the loss of the Micius double-downlink where the link loss was 62 dB.
-[2, 28].
+Through system-level optimization, the source enabled polarization entanglement distribution over 155 km of deployed fiber (66 dB loss).
+The observed brightness and fidelity of the nanophotonic chip source is competitive when compared against work based on bulk crystal sources 18, 19, 20.
+This work marks a milestone in the capability of silicon nanophotonic chips for enabling long range quantum networks.
+One important challenge in distributing polarization-entangled photons over long fibers is PMD, which can degrade entanglement quality.
+In addition to recently proposed compensation techniques 16, 17, microring resonators (MRRs) offer an alternative approach by generating narrowband photon pairs that are inherently less susceptible to PMD.
+The system level optimization concepts proposed by this paper can be applied to other integrated photonics platforms like AlGaAs 42, 43, 36, 44, TFLN 45, 46, SiC 47, 48, USRN 49, AlN 50.
+This source is also useful for satellite to ground experiments because it exceeds the loss of the Micius double-downlink where the link loss was 62 dB 2, 51.
 There are a number of avenues for further integration and improvement.
-For example, many of the filters are based on fiber components and further integration [29, 30, 31] is possible in this aspect enhancing the stability and scalability of the setup.
-We note also that the nonlocal dispersion compensation is performed using a dispersion compensated module that has a significant transmission loss, which should be reduced with further engineering [32, 33].
-We note also that there is a huge potential for utilising dense-wavelength division multiplexing with the silicon nanophotonic chip.
-This will expand the number of channels accessible.
-[23, 34, 22] Finally, other degrees of freedom, such as the time-bins could be added to the setup.
-[35, 36, 37] Taken together, these modifications will dramatically increase the capacity for quantum information transfer within the network.
-4 Methods
-4.1 Chip design and coupling strategy
-The chip involved in the experiment is a 8 mm long silicon waveguide with cross sectional dimension 650 nm in width and 250 nm in height.
-The whole chip is covered by thick SiO2 cladding layer.
-On the chip edges, two-stage inverse tapers are fabricated to expand the chip mode to reduce the mode mismatch with that of the UHNA7 fiber.
-The L0 section acts as buffer region to protect the edge coupler during dicing process, while the length of L1 and L2 tapers are and respectively.
-Short section of UHNA7 fiber (less than 1 cm) is spliced to SMF-28 fiber to assist light collection.
-The loss is optimized to 0.03 dB per splicing junction by performing an in-situ monitoring of the loss during splicing process.
-[16] The measured coupling loss is 0.64 dB per facet and has nice long term stability of dB for more than ten days without using any adhesive.
-4.2 Photon pair wavelength selection
+For example, many of the filters are based on fiber components and further integration 52, 53, 54 is possible in this aspect enhancing the stability and scalability of the setup.
+The nonlocal dispersion compensation is performed using a dispersion compensated module that has a significant transmission loss, which should be reduced with further engineering 55, 56.
+Additionally, there is a huge potential for utilising dense-wavelength division multiplexing with the silicon nanophotonic chip.
+This will expand the number of channels accessible 22, 57, 30.
+Finally, other degrees of freedom, such as the time-bins 58, 59, 60, 61, 62 could be added to the setup to enable high-dimensional entanglement distribution 24, 63, 64.
+Taken together, these strategies can significantly enhance the quantum information capacity of integrated photonic platforms and accelerate the development of scalable, long-range quantum networks.
+We note that a preprint of a related study 65 was posted in March 2025, six months after posting the current work on arXiv.
+METHODS
+Photon pair wavelength selection
 In the SFWM process, two pump photons are absorbed simultaneously, elevating the system to a virtual energy level, followed by the emission of a signal and idler photon to conserve both energy and momentum.
@@
 As illustrated in Fig.
-5.(a), the raw photon coincidence rate was highest when the signal and idler wavelengths were closest to the pump wavelength, indicating that phase matching was most efficient at shorter detunings from the pump.
+4.(a), the raw photon coincidence rate was highest when the signal and idler wavelengths were closest to the pump wavelength, indicating that phase matching was most efficient at shorter detunings from the pump.
 However, this increased photon pair generation does not directly translate into a higher CAR, as shown in Fig.
-5.(b), where the CAR peaks approximately 10 nm away from the pump wavelength.
+4.(b), where the CAR peaks approximately 10 nm away from the pump wavelength.
 This discrepancy can be explained by the balance between photon pair generation and system noise.
+The error bars in Fig.
+4(a,b) represent uncertainties estimated from Poisson statistics and DWDM connector insertion loss variations.
 While photon pairs are generated more efficiently near the pump wavelength, the proximity of the signal and idler to the pump results in less effective rejection of residual pump photons as shown in Fig.
-5.(c).
+4.(c).
 These residual pump photons contribute to accidental coincidences, lowering the CAR.
@@
 The observed peak in CAR at a 10 nm separation from the pump wavelength results from the optimal trade-off between reducing system noise and maintaining sufficient photon pair generation.
-4.3 Methods to improve the fidelity
+Chip design and coupling strategy
+The chip involved in the experiment is a 8 mm long silicon waveguide with cross sectional dimension 650 nm in width and 250 nm in height.
+The whole chip is covered by thick SiO2 cladding layer.
+On the chip edges, two-stage inverse tapers are fabricated to expand the chip mode to reduce the mode mismatch with that of the UHNA7 fiber as shown in Fig.
+5.
+The L0 section acts as buffer region to protect the edge coupler during dicing process, while the length of L1 and L2 tapers are and respectively.
+Short section of UHNA7 fiber (less than 1 cm) is spliced to SMF-28 fiber to assist light collection.
+The loss is optimized to 0.03 dB per splicing junction by performing an in-situ monitoring of the loss during splicing process 25, 66.
+The measured coupling loss is 0.64 dB per facet and has nice long term stability of dB for more than ten days without using any adhesive.
+Low coupling loss is critical for practical entangled photon-pair sources, not only because it enhances the overall brightness, but also because it significantly impacts the CAR 67, as shown in Fig.
+6.(a).
+The relationship between the CAR, the on-chip raw coincidence , and the system efficiencies can be expressed as:
+In SFWM, the pair generation rate scales quadratically with pump power, while the background noise (Raman scattering and unfiltered pump) scales linearly .
+For a fixed detected coincidence rate , it follows that , , and hence .
+The denominator of Eq.
+2 thus scales as
+indicating that CAR increases monotonically with coupling efficiency .
+Nonlocal dispersion compensation
 In this experiment the deployed fiber is not dark and contains broadband noise within the optical channel of interest.
 This noise lowers the CAR value affecting the fidelity of the observed quantum state.
-The relationship between fidelity and CAR is given by .
+The maximum achievable fidelity is limited by the CAR and is given by .
 A 0.8 nm filter is installed before the SNSPDs, and the dark count noises are effectively removed from around 4 Mcps to less than 200 cps.
-The CAR is increased from 2 to 15, which corresponds to a theoretical fidelity improvement from 67% to 94%.
-Additionally, chromatic dispersion presents a significant challenge to achieving high entanglement fidelity.
+As shown in Fig.
+6.(b), the CAR is increased from 2 to 15, which corresponds to a fidelity upper bound improvement from 67% to 94%.
+Additionally, chromatic dispersion presents a significant challenge to achieve high entanglement fidelity.
 The dispersion value at 1539 nm is around ps/(nmkm) in the fiber.
@@
 The total dispersion compensation provided by the DCM at 1561 nm is around -1360 ps/nm, as data provided by the manufacturer.
-Effective dispersion compensation raises the CAR from 15 to 91, corresponding to an increase in theoretical fidelity from 94% to 99%.
+As shown in Fig.
+6.(b), effective dispersion compensation raises the CAR from 15 to 91, corresponding to an increase in theoretical fidelity from 94% to 99%.
 However, the measured fidelity in practice is lower than this theoretical value, primarily due to the effects of PMD.
 It is important to note that the DCM introduces 4.3 dB insertion loss, which does impact the coincidence rate.
-This DCM-related loss may be mitigated in future work where the waveguide platform could be designed to generate photons with narrower wavelength bandwidth.
-[38, 12, 22, 17, 39]
-4.4 System calibration
+This DCM-related loss may be mitigated in future work where the waveguide platform could be designed to generate photons with narrower wavelength bandwidth 68, 26, 30, 31, 69.
+Phase stabilization in deployed fibers
+In the 93 km asymmetric configuration, phase fluctuations are more pronounced because only one photon passes through the long fiber while the other is detected locally.
+In this case, environmental changes such as temperature-induced fiber expansion or contraction directly translate into phase drift between the two optical paths.
+In contrast, the near-symmetric entanglement distribution experiment utilizes both the 93 km and 62 km deployed fiber segments, with each photon of the entangled pair traversing one segment.
+This configuration benefits from partial passive phase noise cancellation: since both fibers are deployed underground in similar metropolitan conditions, environmental fluctuations tend to affect both arms similarly, leading to correlated phase changes and reduced differential phase drift.
+The phase stability experiments were performed on the 62 km fiber segment, and the results are shown in Fig.
+6.(c).
+The observed delay drift in this configuration is representative of the deployed metropolitan fiber environment.
+Every 5 minutes, the Sagnac pump splitting ratio is switched from 1:1 to 1:0 by rotating the HWP in the pump polarization control unit in Fig.
+1.(a), and the waveplates in the state analyzer in Fig.
+1.(e) are adjusted to measure the HH coincidence histogram.
+This temporary configuration doubles the pump power and hereby increases the coincidence count by a factor of four, enhancing the precision of delay estimation.
+The delay is extracted by fitting the coincidence histogram with a Gaussian envelope and summing five bins around the peak (when using 50 ps coincidence window), with the central bin identified as the reference delay.
+During the subsequent 5-minute measurement period, the pump splitting ratio is reverted to 1:1, and the previously extracted delay value is applied.
+This calibration cycle is repeated at 5-minute intervals.
+As shown in Fig.
+6.(c), the uncompensated phase drift can exceed over time.
+A 60-minute feedback interval provides moderate suppression, whereas a 5-minute compensation cycle confines the delay variation within for more than 70 hours, achieving a stabilization accuracy of 99.6%.
+These results highlight the importance of employing high-brightness entangled photon-pair sources.
+Without sufficient brightness, longer integration times are required to obtain statistically significant data, during which uncorrected phase drift can significantly degrade the CAR and reduce the fidelity of the reconstructed quantum state.
+System calibration
 The polarization shift caused by birefringence is compensated by the QWP and HWP in the measurement setup.
@@
 This includes all the fiber based and free space optical components in the experiment.
-4.5 Single photon detection and coincidence counting
-In the 93 km entanglement distribution experiment, the single photons are detected with SNSPDs, with the efficiency for idler channel at 45% while efficiency for signal channel is at 78%.
-Single photon count rate is 4.8 kcps for signal arm and 5.6 Mcps for idler arm.
-The delay cause by the deployed fiber is 457,369,970 ps which is equal to the distance measured by OTDR.
-The coincidence window used in local measurement is 200 ps while with deployed fiber is 60 ps for accidental suppression and higher CAR.
-The narrow does result in a lower photon pair rate.
+Single photon detection and coincidence counting
+The entangled photons are detected by SNSPDs.
+The detection efficiency of the SNSPDs is tunable via the device bias current: higher bias increases both detection efficiency and dark count rate, but can reduce the maximum detectable count rate by making the detector more susceptible to latching.
+For source characterization measurements, the SNSPD efficiency was intentionally set to a low value (45% and 49%) to increase the pump power substantially without saturating the detectors.
+In the 93 km entanglement distribution experiment, the single photons were detected with SNSPDs set to higher efficiency: 78% for the signal channel and 45% for the idler channel.
+The single-photon count rates were 4.8 kcps for the signal arm and 5.6 Mcps for the idler arm.
+The delay caused by the deployed fiber is 457,369,970 ps, matching the distance measured by OTDR.
+The coincidence window used in local measurements is 200 ps, while with deployed fiber is 60 ps for accidental suppression and higher CAR.
+The narrow results in a lower photon-pair rate.
 Reducing the background noise (detector dark counts, Raman noise in the source, and stray light coupling into the fiber network) should further improve the CAR and fidelity.
-When the deployed fiber distance is extended to 155 km, the singles count rate on signal and idler arm are around 15 kcps and 40 kcps respectively.
-Although this low single count rate will not saturate the detector, the SNSPD are set to 56% efficiency to reduce the dark count to 10 cps.
-The optical delay caused by the 62 km fiber is 302,113,173 ps which is equal to the length characterization by OTDR.
-4.6 Fiber link information
-A measurement of polarization stability of deployed fibers over a continuous period of 8 hours is conducted and the polarization state rotates (93 km link) and (62 km link) on the Poincare sphere.
+When the deployed fiber distance was extended to 155 km, the singles count rates on the signal and idler arms were around 15 kcps and 40 kcps, respectively.
+Although this low single-photon count rate does not saturate the detector, the SNSPDs were set to 56% efficiency by lowering the bias current to reduce the dark count rate to 10 cps.
+The optical delay caused by the 62 km fiber is 302,113,173 ps, consistent with the length characterization by OTDR.
+Fiber link information
+A measurement of polarization stability of deployed fibers over a continuous period of 8 hours is conducted and the polarization state rotates (93 km link) and (62 km link) on the Poincaré sphere.
 The results in Fig.
@@
 After inserting the filter, the background noise in both channels is reduced to less than 200 cps.
-Acknowledgements
+RESOURCE AVAILABILITY
+Lead contact
+Requests for further information and resources should be directed to and will be fulfilled by the lead contact Alexander Ling (alexander.ling@nus.edu.sg).
+Materials Availability
+This study did not generate new unique materials.
+Data and code availability
+The data reported in this study are available from the lead author upon reasonable request.
+This paper does not report any original code.
+Any additional information required to reanalyze the data reported in this paper is available from the lead contact upon request.
+ACKNOWLEDGMENTS
 This research is supported by the National Research Foundation, Singapore, and A*STAR under its CQT Bridging Grant and Quantum Engineering Programme (Award No.
@@
 We acknowledge funding support from the National Research Foundation, Singapore and A*STAR under its Quantum Engineering Programme (National Quantum-Safe Network, NRF2021-QEP2-04-P01) and Netlink Trust for the provisioning of the fiber network.
-The authors acknowledge the assistance of setup alignment and literature review from Arya Chowdhury, and Lim En Teng, and consultation from Dr.
-Tanvirul Islam.
-Declarations
-The authors declare no conflicts of interest.
+The authors acknowledge the assistance of setup alignment and literature review from Arya Chowdhury, and En Teng Lim, and consultation from Dr.
+Tanvirul Islam, Dr.
+Haoran Zhang, and Dr.
+Shihao Ru.
+AUTHOR CONTRIBUTIONS
+J.D., X.Z., D.T.H.T., and A.L.
+conceived the experiment; G.F.R.C.
+and H.G.
+contributed to materials processing and fabrication; J.D.
+performed chip characterization; J.D.
+and X.Z.
+conducted entanglement measurements and analyzed the data; J.D., D.T.H.T., and A.L.
+wrote the manuscript; A.L.
+and D.T.H.T.
+supervised the project and acquired funding.
+DECLARATION OF INTERESTS
+Information related to the edge coupler design used in this work is covered by a pending international patent application (Application No.
+PCT/SG2024/050518, titled “Coupling Platform for Photonic Chips”) filed by Jinyi Du, Dawn T.
+H.
+Tan, and Alexander Ling.